Use Cramer's Rule to solve the system of linear equations.

xa + 2xc = 6
-3xa + 4xb + 6xc = 30
-xa - 2xb + 3xc = 8
1
Expert's answer
2021-09-03T04:59:19-0400

"\begin{gathered}
\Delta=\left|\begin{array}{ccc}
x & 0 & 2 x \\
-3 x & 4 x & 6 x \\
-x & -2 x & 3 x
\end{array}\right|=x\left(12 x^{2}+12 x^{2}\right)-0+2 x\left(6 x^{2}+4 x^{2}\right) \\
=24 x^{3}+20 x^{3}=44 x^{3} \\
\Delta_{1}=\left|\begin{array}{ccc}
6 & 0 & 2 x \\
30 & 4 x & 6 x \\
8 & -2 x & 3 x
\end{array}\right|=6\left(12 x^{2}+12 x^{2}\right)-0+2 x(-60 x-32 x) \\
=144 x^{2}-184 x=4 x(36 x-46) 
\\\Delta_{2}=\left|\begin{array}{ccc}
x & 6 & 2 x \\
-3 x & 30 & 6 x \\
-x & 8 & 3 x
\end{array}\right| \\
=x(90 x-48 x)-6\left(-9 x^{2}+6 x^{2}\right)+2 x(-24 x+30 x) \\
=42 x+18 x^{2}+12 x^{2}=42 x+30 x^{2}=2 x(21+15 x) \\
\Delta_{3}=\left|\begin{array}{ccc}
x & 0 & 6 \\
-3 x & 4 x & 30 \\
-x & -2 x & 8
\end{array}\right|=x(32 x+60 x)-0+6\left(6 x^{2}+4 x^{2}\right) \\
=92 x^{2}+60 x^{2}=152 x^{2}
\end{gathered}"

"Now, a=\frac{\Delta_{1}}{\Delta}=\frac{4 x(36 x-46)}{44 x^{3}}=\frac{(36 x-46)}{11 x^{2}},\\b=\frac{\Delta_{2}}{\Delta}=\frac{2 x(21+15 x)}{44 x^{3}}=

\frac{(21+15 x)}{22 x^{2}},\\ c=\frac{\Delta_{3}}{\Delta}=\frac{152 x^{2}}{44 x^{3}}=\frac{38}{11 x}"